OpenStudy (anonymous):

I need help plz i do not get this at all What is the kinetic energy, in J, of: (A) an Ar atom moving with a speed of 650m/s? (B) a mole of Ar atoms moving with a speed of 650m/s (Hint: 1 amu =1.66 * 10^-27 kg)?

OpenStudy (anonymous):

I need help with finding the mass of the Ar atom plz and thank you

OpenStudy (aaronq):

use the formula: Kinetic Energy = \(mv^2\)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Now I have a question about Hess's Law MnO2 + CO MnO + CO2 H = -151 kJ Mn3O4 + CO 3MnO + CO2 H = -54 kJ 3Mn2O3 + CO 2Mn3O4 + CO2 H = -142 kJ A) Use the above thermochemical equations to determine H for the following: 2MnO2 + CO Mn2O3+ CO2

OpenStudy (anonymous):

I know how to do it but I can't figure which equation to flip